solve for x and y in the following problem
18 x +y = 26, 6x -3y = 24

Answer by checkley71(8403)

18X+Y=26 & 6X-3Y=24 MULT EQ #2 BY -3 AND THEN ADD THESE TQO EQ WE GET
18X+Y=26
-18X+9Y=-72
------------
10Y=-46 Y=-46/10 OR Y=-4.6 THEN 18X-4.6=26 OR 18X=30.6 OR X=30.6/18 OR X=1.7